Find area of the region outside the circle r=2 and inside the lemniscate r^2=8cos2α
1
Expert's answer
2019-04-29T11:46:54-0400

Find angles, when the circle inside the lemniscate:


"r^2 = 8 \cos(2\alpha) \geq 2^2""\cos(2\alpha)\geq 1/2""\frac{1}{6} \left(6 \pi n-\pi \right)<\alpha <\frac{1}{6} \left(6 \pi n+\pi \right),\,\,n \in \mathbb{Z}"

Using periodicity over "\alpha" , lets convert usual convension over domain "0<\alpha<2\pi" , to more convenient in this case,


"-\pi/2<\alpha<3\pi/2"

This correspond to n = 0 and n = 1.

Finally, domain of angles, when the circle inside the lemniscate:


"-\frac{\pi}{6} <\alpha <\frac{\pi}{6},\,\,\,\,\text{and}\,\,\,\,\,\frac{5\pi}{6} <\alpha <\frac{7\pi}{6}"

AREA OF A REGION BOUNDED BY A POLAR CURVE:


"S_0 = \frac12 \int_{-\pi/6}^{\pi/6} (8cos(2α)-2^2)d\alpha = 2(\sin(2α)-\alpha)|_{-\pi/6}^{\pi/6}=2 \sqrt{3}-\frac{2 \pi }{3}"

As integration in region "\frac{5\pi}{6} <\alpha <\frac{7\pi}{6}" would give the same result "S_1 = S_0" ,

therefore final result - area of the region outside the circle and inside the lemniscate equals to


"4 \sqrt{3}-\frac{4 \pi }{3}"